# Using matrix-product states for time-series machine learning

### v1

@@ V Discussion and Conclusion
Here we have presented MPSTime, an end-to-end framework for training and applying matrix-product states to time-series machine-learning problems. We show that a generative model parameterized by an MPS can effectively capture complex, high-dimensional joint time-series distributions, making it particularly well-suited to handling the challenges of time-series data. Leveraging the joint distribution encoded by the MPS, we demonstrate the versatility of our model by tackling important problems in time-series ML – imputation and classification – within a flexible and interpretable statistical framework. To our knowledge, this is the first framework for inferring joint distributions directly from time-series data. This provides a unified approach for tackling important problems, including classification, imputation, and synthetic data generation. Previously, existing approaches in the ML literature have treated these tasks in a disjoint manner, lacking a clear and common conceptual basis. Further, for the applications in this work, our MPS-based generative model only requires moderate bond dimension $\chi_{\rm max}$ and physical dimension $d$, striking a compelling balance between model expressiveness and computational efficiency.
In conclusion, we first validated our method on a range of synthetically generated datasets, showcasing its remarkable ability to recover missing data under diverse and challenging conditions. Our results underscore the model’s capacity to impute values in datasets with complex temporal correlations, achieving near-ground-truth accuracy even when faced with noisy and limited observations. We further benchmarked the method on three real-world datasets, where it demonstrated competitive and often superior performance relative to several classical and state-of-the-art imputation algorithms. Having demonstrated the promising imputation capabilities of the MPS, we then showcased its flexibility to adapt to classification tasks on the same real-world datasets. The MPS classifier exhibited robust performance across all three datasets, achieving mean classification accuracies comparable to those of state-of-the-art classification algorithms.
@@ Acknowledgements.
This research was supported by the University of Sydney School of Physics Foundation’s Grand Challenge initiative. S.M. acknowledges support from the Australian Research Council (ARC) via the Future Fellowship, ‘Emergent many-body phenomena in engineered quantum optical systems’, project no. FT200100844.
@@ Appendix A Data pre-processing details
Here we discuss the details of the data transformations we applied to each raw time-series dataset prior to encoding them with the feature map. For a given time-series dataset of $N$ instances and $T$ data points (samples) per instance, represented by an $N\times T$ data matrix $\bm{X}$, the outlier-robust sigmoid transformation is given by [36]:
@@
where $\bm{X^{\prime}}$ is the normalized time-series data matrix, $\bm{X}$ is the un-normalized time-series data matrix, $m_{\bm{X}}$ is the median of $\bm{X}$ and $r_{\bm{X}}$ is its interquartile range. In the case of a scaled robust sigmoid transform, an additional linear transformation, commonly referred to as MinMax, is applied to $\bm{X^{\prime}}$ such to rescale it to the target range $[a,b]$:
@@
where $\bm{X^{\prime\prime}}$ is the scaled robust-sigmoid transformed data matrix, $x^{\prime}_{\text{min}}$ and $x^{\prime}_{\text{max}}$ are the minimum and maximum of $\bm{X^{\prime}}$, respectively. In cases where we only apply the linear rescaling, we use the same transformation in Eq. 26, but applied to the raw data matrix $\bm{X}$ rather than the sigmoid transformed data matrix $\bm{X^{\prime}}$.
When we evaluated a trained model on unseen data, the unseen data was preprocessed with the values $m_{\bm{X}}$, $r_{\bm{X}}$, $x^{\prime}_{\text{min}}$, and $x^{\prime}_{\text{max}}$ extracted from the training dataset. An undesirable implication of this is that occasionally test time series were preprocessed to map outside of the range $[a,b]$. For each of these time-series instances:
@@
    If the minimum was less than $a$, the time series was shifted up so it had a minimum value of $a$.
@@
    After that, if the maximum was larger than $b$, the time series was rescaled so it was at most $b$.
These outliers had no meaningful bearing on the results. For example, for the default UCR train/test split of the ECG dataset, only $5$ out of the $100$ time series in the test set were shifted up by $0.1\%$, $0.1\%$, $0.2\%$, $1\%$, and $1.7\%$, respectively. No time series were rescaled. For the default UCR train/test split of the Power Demand dataset, only $34$ out of the $1029$ test time series were altered: $13$ were shifted up, the five largest shifts were: $1.9\%$, $2.1\%$, $2.1\%$, $4.4\%$ and $5.1\%$. A total of $23$ time series were rescaled, with the five largest rescales being by $2.6\%$, $3.0\%$, $4.1\%$, $4.5\%$ and $8.8\%$.
When evaluating imputation performance in Sec. III.2, the mean absolute error (Eq. (23)) was computed on the raw (non-preprocessed) data.
@@ Appendix B Impact of artificial broadening from finite basis representation on conditional imputation
In this work, we train an MPS to encode a probability distribution of time-series amplitudes. As these amplitudes are originally real-valued and continuous ($x_{t}\in\mathcal{R}$), they must first be mapped to discrete quantum states (|$\ket{x_{t}}$) to be compatible with the MPS framework. To achieve this discretization, we project the amplitudes onto a truncated basis set of $d$ functions, $\phi_{t}(x_{t})=\{b_{1}(x_{t}),b_{2}(x_{t}),\ldots,b_{d}(x_{t})\}_{t}$. Here, the choice of $d$ determines the resolution of the discretization, with larger $d$ providing a finer approximation of the original continuous time-series amplitudes.
Notably, when discretizing with a finite $d$, the reduced density matrix (RDM) at each site can assign artificially inflated probabilities to states that would otherwise be unlikely under the true distribution. This results in an artificial broadening of the encoded probability distributions, as depicted in Fig. 9. In the context of imputation, such broadening can lead to errors, as randomly sampling from the distribution may include unrepresentative states, disproportionately skewing quantities such as the expectation.
Figure 9: Distribution broadening due to the finite basis representation of a continuous value. Given a continuous value in the compact domain $x^{\prime}\in[-1,1]$, e.g., $x^{\prime}=0.3$, represented by the dotted black line, the corresponding conditional distribution under the finite Legendre basis representation is shown for varying $d$. Summary statistics (mean, median and mode) derived from each distribution are shown as colored markers.
For the purposes of imputation, however, the goal is to obtain a single representative estimate from the conditional distribution at a particular time point. To this end, we chose to investigate deterministic summary statistics derived from the conditional distribution, as opposed to individual random samples. We then defined an ‘encoding error’ $\epsilon_{d}$ to quantify the discrepancy introduced by the finite basis approximation:
@@
where $x_{i}$ is a continuous value in the encoding domain, and $\hat{x}_{i,d}$ is a summary statistic derived from the conditional probability distribution at site $i$, after encoding $x_{i}$ as a discrete quantum state (via the encoding $\phi_{i}(x_{i})$) using a finite basis with dimension $d$. In Fig. 10, we compare the encoding error for three summary statistics – the expectation (mean), median, and mode – as a function of $d$ and the continuous value in the Legendre encoding domain $x_{i}\in[-1,1])$. As expected, the mean, shown in Fig. 10(a), exhibits large encoding errors across the domain, particularly for smaller $d$, as a result of its sensitivity to the artificial broadening of the distribution. The mode (Fig. 10(b)), while competitive for higher $d$, exhibits significant error spikes near the domain boundaries which can lead to erroneous imputations. The median, shown in Fig. 10(c), was found to consistently outperform other summary statistics, providing the lowest average error across the encoding domain. Motivated by these empirical findings, we chose to adopt the median, given that it best mitigates the effects of artificial broadening and ensures the most representative single point estimate from the encoded distribution.
@@
Figure 10: Encoding error over the Legendre domain. The absolute error between a known value $x^{\prime}$ in the encoding domain, and a summary statistic for the probability distribution corresponding to the conditional density matrix $\rho=\phi^{\dagger}(x^{\prime})\phi(x^{\prime})$ (note that the Legendre feature map $\phi$ is time independent), which we refer to as the encoding error $\epsilon_{d}$ (as in Eq. (27)) is shown for (a) the mean (expectation); (b) the mode; and, (c) the median. In subplot (i) of each main panel, the encoding error when using the summary statistic as the single point estimator, is shown for varying $d$ and $x^{\prime}$ across the compact Legendre encoding domain $x^{\prime}\in[-1,1]$. In each subplot (ii), we show the encoding error for a few selected values of $d$.
Appendix C Hyperparameter tuning for classification
@@
For classification, the physical dimension $d$, maximum bond dimension $\chi_{\rm max}$ and learning rate $\eta$ were determined by hyperparameter tuning. The particular hyperparameter ranges we searched for each empirical dataset are summarized in the table below:
@@
$d$ [2, 15] [2, 4] 3
$\eta$ [0.001, 10.0] [0.001, 10.0] 1.0
$\chi_{\rm max}$ [15, 50] [15, 30] 30
encoding Legendre Legendre Fourier
To carry out hyperparameter tuning for ECG and Power Demand, we used an off-the-shelf Julia implementation of the adaptive particle swarm optimization algorithm [66] with stratified 5-fold cross-validation. Since the Astronomy dataset was derived as a subset of the original KeplerLightCurves UCR dataset, no published results exist for the baseline methods. To ensure a fair comparison in this context, we evaluated the baseline methods using their default configurations, i.e., their out-of-box performance. While these default parameters are typically based on extensive hyperparameter studies, such investigations are not the focus of the present work. Instead, for our MPS-based classifier, we selected a parameter set that performed well in our preliminary experiments. It is important to note that this choice of parameters does not confer an unfair advantage: our 30-fold resampling strategy ensures that any given configuration performing well on one fold may not necessarily do so in others, as is the case for the default parameters of the baseline methods.

### v2

@@ V Discussion and Conclusion
In this work we have introduced MPSTime, a flexible algorithmic framework for learning complex time-series distributions using matrix-product states (MPS) that can be applied to a range of time-series machine-learning (ML) problems, including classification and imputation. This capacity to tackle diverse time-series ML tasks within a unified probabilistic framework sets MPSTime apart from many conventional ML algorithms which focus on a single problem class in isolation (e.g., disjoint algorithms for time-series classification, imputation, etc.). In addition, the ability to extract the learned temporal dependencies directly from the MPS (via the reduced density matrix) grants MPSTime a level of interpretability that is lacking in many other time-series ML approaches. For the machine learning applications presented here, MPSTime required only moderate bond dimension ($\chi_{\rm max}=20-160$) and physical dimension ($d=2-15$), achieving a favorable balance between expressiveness and computational cost. Taken together, our results support the MPS ansatz as an efficient and useful way to distill complex temporal dependencies contained in time-series data, paving the way for future work to extend our approach to tackle a wider range of time-series problems, including forecasting and synthetic data generation.
Building on a growing body of work applying MPS to machine learning, this work contributes several significant innovations to an emerging field. For the first time, we develop an end-to-end framework for training MPS to approximate the joint distribution of continuous-valued sequential (time-series) data. While generalized feature maps (e.g., Legendre and Fourier basis) have been used to encode continuous-valued inputs of at most four dimensions (i.e., an MPS of four sites) [20], our approach scales to continuous-valued time-series data of up to 100 samples (i.e., an MPS of 100 sites) and beyond, while remaining practical to run, even on real-world datasets. We introduce two complementary error reduction techniques – leveraging the conditional median as a robust point estimate for time-series imputation, and a rejection sampling scheme that curbs the propagation of encoding-related errors with site-wise sampling. Together, our methods overcome the inherent limitations of discrete encodings for continuous-valued time-series data, enabling accurate imputation and robust data generation on long sequences. We anticipate that such algorithmic advancements will provide a foundation for MPS-based generative modeling applications. We have made our implementations of the imputation, classification, MPS-analysis, and hyperparameter optimization algorithms used in this work publicly available in the comprehensively documented code package MPSTime. Our implementation has been extensively optimized, and, unlike many publicly available time-series ML algorithms, does not require specialized knowledge or hardware to run.
While our approach achieves performance comparable to state-of-the-art benchmarks on a range of time-series ML tasks, there is room for further improvement. First, as a preliminary investigation, we focused on a small number of representative datasets to demonstrate the viability of MPS-based algorithms for time-series ML. While a comprehensive benchmark falls outside the scope of this study, it offers a clear road map for future investigations. A limitation of the fixed-length MPS geometry is that the current implementation of MPSTime can only handle time-series instances of fixed length. Although this can be partially addressed through time-series windowing, such an approach neglects potentially crucial long-range temporal dependencies that extend beyond the chosen window length. A promising avenue for future work is to explore translationally invariant MPS (TI-MPS) [76] which have long been used in quantum many-body physics to approximate one-dimensional systems in the thermodynamic limit [77]. Owing to their flexible architecture, such TI-MPS can be naturally adapted to time series of varying lengths, opening up the possibility of time-series forecasting across arbitrary horizons. Furthermore, for stationary time series, where the statistical properties are invariant to time shifts [78], incorporating translational invariance into the MPS geometry could enhance out-of-distribution generalization. As found in Sec. IV.3, the behavior of MPSTime can be approximately thought of as a nearest-neighbor-like efficient compression of the training distribution, which is desirable in many settings, but limits its out-of-distribution generalization. One direction in which this functionality could be incorporated is through TI-MPS, which should generalize a stationary correlation structure to new time series; in the simple test case of noisy sinusoids with variable phase studied in Sec. IV.3, a TI-MPS should be able to accurately impute missing values in sinusoids with arbitrary phases after training on examples of only a single phase. This can have particular advantages for very large time-series datasets where nearest neighbor can become prohibitively slow.
There are many promising extensions of MPSTime to problem classes beyond classification and imputation, demonstrated here, including to anomaly detection [79], regression (of extrinsic variables from time series [80]), forecasting, and synthetic data generation. Such extensions are made relatively straightforward as they can be developed from a common underlying probabilistic model, MPSTime, similarly to other flexible learning algorithms for time-series modeling [38]. The application to synthetic time-series generation is a particularly promising direction given our results presented in Sec. IV.1, where conditionally sampling from a trained MPS can yield an ensemble of time series that closely resemble the training dataset. Future work should also assess the performance of MPSTime relative to a range of other flexible generative models for time-series data, including recent developments in this space using techniques including variational autoencoders (VAEs), generative adversarial networks (GANs), and denoising diffusion probabilistic models [38].
In conclusion, here we have demonstrated the usefulness of the MPS ansatz in efficiently capturing complex time-series data structures and introduced an open and extendable algorithmic implementation, MPSTime, that implements the method. The high performance of MPSTime in classification and imputation settings along with its potential for interpretability make it a promising and flexible new approach to modeling complex time-series data, providing a common theoretical foundation on which algorithms for tackling a range of time-series problem classes can be developed.
@@ Acknowledgements.
This research was supported by the University of Sydney School of Physics Foundation’s Grand Challenge initiative. S.M. acknowledges support from the Australian Research Council (ARC) via the Future Fellowship, ‘Emergent many-body phenomena in engineered quantum optical systems’, project no. FT200100844. B.D.F. acknowledges support from the Australian Research Council (FT240100418).
@@ Appendix A Data pre-processing details
Here we discuss the details of the data transformations we applied to each raw time-series dataset prior to encoding them with the feature map in Eq. (3). For a given time-series dataset of $N$ instances and $T$ data points (samples) per instance, represented by an $N\times T$ data matrix $\bm{X}$, we used an outlier-robust sigmoid transformation [39]:
@@
where $\bm{X^{\prime}}$ is the normalized time-series data matrix, $\bm{X}$ is the un-normalized time-series data matrix, $m_{\bm{X}}$ is the median of $\bm{X}$ and $r_{\bm{X}}$ is its interquartile range (IQR). In the case of a scaled robust sigmoid transform, an additional ‘min–max’ linear transformation is applied to $\bm{X^{\prime}}$ such to rescale it to the target range $[a,b]$:
@@
where $\bm{X^{\prime\prime}}$ is the scaled robust-sigmoid transformed data matrix, $x^{\prime}_{\text{min}}$ and $x^{\prime}_{\text{max}}$ are the minimum and maximum of $\bm{X^{\prime}}$, respectively. As discussed in Sec. II.1 for time-series imputation we only used the linear transformation in Eq. 26, but applied to the raw data matrix $\bm{X}$ rather than the sigmoid-transformed data matrix $\bm{X^{\prime}}$. For classification, we used the scaled robust sigmoid transformation as defined in Eq. (26).
When we evaluated a trained model on unseen data, the unseen data was normalized using the values $m_{\bm{X}}$, $r_{\bm{X}}$, $x^{\prime}_{\text{min}}$, and $x^{\prime}_{\text{max}}$ estimated from the training dataset to avoid information leakage between the training and test sets. An undesirable implication of this necessary processing is that occasionally some test time series were mapped outside of the range $[a,b]$. For each of these time-series instances:
@@
    If the minimum was less than $a$, the time series was shifted up so that it had a minimum value of $a$.
@@
    After shifting the time series, if the new maximum was larger than $b$, the time series was linearly rescaled so that the maximum was at most $b$.
These outliers had no meaningful bearing on classification performance. For example, for the default UCR train/test split of the ECG dataset, only $5$ out of the $100$ sigmoid-transformed time series in the test set were shifted up by $0.1\%$, $0.1\%$, $0.2\%$, $1\%$, and $1.7\%$, respectively. No time series were rescaled due to maximums exceeding the encoding domain after shifting. For the default UCR train/test split of the Power Demand dataset, only $34$ out of the $1029$ sigmoid-transformed test time series were altered. Specifically, $13$ time-series instances were shifted up, the largest of which were $1.9\%$, $2.1\%$, $2.1\%$, $4.4\%$ and $5.1\%$. A total of $23$ time series were rescaled, with the five largest rescales being by $2.6\%$, $3.0\%$, $4.1\%$, $4.5\%$ and $8.8\%$.
When evaluating imputation performance in Sec. III.2, the mean absolute error (MAE) defined in Eq. (23) was computed on the raw (non-processed) data to ensure errors were presented in meaningful and comparable units across all imputation algorithms.
@@ Appendix B Impact of artificial broadening from finite basis representation on conditional imputation
In this work, we train an MPS to encode a probability distribution of time-series amplitudes. As these amplitudes are real-valued and continuous, with each $x_{t}\in\mathbb{R}$, they must first be mapped to discrete quantum states, |$\ket{x_{t}}$, to be compatible with the MPS framework. To achieve this discretization, we project the amplitudes onto a truncated basis set of $d$ functions, $\phi_{t}(x_{t})=\{b_{1}(x_{t}),b_{2}(x_{t}),\ldots,b_{d}(x_{t})\}_{t}$. Here, the choice of $d$ determines the resolution of the discretization, with larger $d$ enabling a finer approximation of the original continuous time-series amplitudes.
As a result of this discretization, the reduced density matrix (RDM) at each site can assign artificially inflated probabilities to states that would otherwise be unlikely under the true data distribution. This effect is depicted visually in Fig. 10, which shows the ‘artificial broadening’ of the conditional distribution, given a ground-truth continuous value, $x^{\prime}=0.3$ (vertical black line), encoded as a discrete state via a Legendre feature map with $d$ basis functions, as in Eq. (3). For low physical dimension $d=4$ (blue), corresponding to a coarser approximation, the distribution is broad and skewed. Repeatedly sampling from this distribution will tend to yield estimates far from $x^{\prime}=0.3$ (vertical black line). As $d$ increases (e.g., $d=12$, red) – yielding a more fine-grained approximation of the continuous value – the mass concentrates around $x^{\prime}=0.3$, and measures of central tendency such as the median (red diamond marker) and mode (red triangle marker) converge toward the true value, with the latter two summary statistics exhibiting the greatest robustness.
Figure 10: Distribution broadening due to the finite basis representation of a continuous value. Given a continuous value in the compact domain $x^{\prime}\in[-1,1]$, e.g., $x^{\prime}=0.3$, represented by the vertical black line, the corresponding conditional distribution under the finite Legendre basis representation is shown for varying MPS physical dimension $d$. Three measures of the central tendency of a distribution: mean (cross marker), median (diamond marker), and mode (triangle marker) are shown for $d=4$ (blue), $d=8$ (gold), $d=10$ (green), and $d=12$ (red).
In this work, the imputation problem we address involves inferring a single point estimate for each missing time-series sample. In this setting, rather than sampling from the conditional distribution, for a sufficiently large $d$, we aimed to use a measure of central tendency as our point estimate. We then defined an ‘encoding error’ $\epsilon_{d}$ to quantify the discrepancy introduced by the finite basis approximation:
@@
where $x_{i}$ is a continuous value in the encoding domain, and $\hat{x}_{i,d}$ is a summary statistic capturing the central tendency of the conditional probability distribution at site $i$ after encoding $x_{i}$ via the feature map $\phi_{i}(x_{i})$ with $d$ basis functions. In Fig. 11, we compare the encoding error for three measures of central tendency – the expectation (mean), median, and mode – as a function of $d$ and uniformly spaced values in the Legendre encoding domain ($x_{i}\in[-1,1]$). As expected, the mean, shown in Fig. 11(a), gives rise to large encoding errors across the domain, particularly for smaller $d$, as a result of its sensitivity to the artificial broadening of the distribution. The mode (in Fig. 11(b)), while competitive for higher $d$, exhibits significant error spikes near the domain boundaries which can lead to erroneous imputations. The median, shown in Fig. 11(c), was found to consistently outperform other summary statistics, providing the lowest average error across the encoding domain. Motivated by these empirical findings, we chose to adopt the median for MPS-based time-series imputation, given that it best mitigates the effects of artificial broadening and ensures a sufficiently representative single point estimate from the encoded distribution.
@@
Figure 11: Encoding error of continuous values over the Legendre domain under a finite basis approximation. The absolute error between a known value $x^{\prime}$ in the encoding domain, and a summary statistic for the probability distribution corresponding to the conditional density matrix $\rho=\phi^{\dagger}(x^{\prime})\phi(x^{\prime})$ (note that the Legendre feature map $\phi$ is time independent), which we refer to as the encoding error $\epsilon_{d}$ (as in Eq. (27)) is shown for (a) the mean (expectation); (b) the mode; and, (c) the median. Top row (i): shows the encoding error, indicated by the color scale, when using a measure of central tendency as a single point estimator for varying $d$ and $x^{\prime}$ across the compact Legendre encoding domain $x^{\prime}\in[-1,1]$. Bottom row (ii): each column – (a)(ii), (b)(ii), (c)(ii) shows the encoding error for selected values of physical dimension, $d=4$ (blue), $d=12$ (green) and $d=16$ (red), corresponding to the expectation error in (a)(i), mode error in (b)(i), and median error in (c)(i), respectively.
Appendix C MPSTime hyperparameter tuning
C.1 Imputation hyperparameter tuning
For our imputation experiments in Sec. III.2, the MPS physical dimension $d$, maximum bond dimension $\chi_{\rm max}$, and learning rate $\eta$ were determined by hyperparamater optimization. For the learning rate $\eta$, we searched in logarithmic space to efficiently cover multiple orders of magnitude. We note that in Sec. III.2.1 $\chi_{\rm max}$ was held fixed, as we sought to investigate the relationship between bond dimension (model expressibility) and imputation accuracy. The number of sweeps was fixed to $n_{\rm sweeps}=10$ across all experiments, and we used a min–max pre-processing transformation for the data, as described in Sec. II.1. To search the space of hyperparameters bounded by the ranges defined in Table. 2 we used an off-the-shelf implementation of Latin Hypercube Sampling (LHS) [81] with 5-fold cross-validation (CV). LHS is a random search variant suited to efficiently sampling from multi-dimensional distributions. For each CV fold, $250$ MPS models (i.e., unique hyperparameter combinations) were evaluated based on their mean MAE across the 5% to 95% data missing range, averaged across all held-out validation instances.
Table 2: MPSTime imputation hyperparameter ranges for each dataset in Sec. III.2.
Hyperparameter NTS1-4 NTS5 §Empirical
$d$ [5, 15] [5, 15] [5, 15]
$\eta$ [0.001, 0.5] [0.001, 0.5] [0.001, 0.5]
$\chi_{\rm max}$ [20, 80]* [20, 160] [20, 40]
NTS1–4 refers to datasets NTS1, NTS2, NTS3, NTS4 described in Sec. III.1.1. *For NTS1, we hold $\chi_{\rm max}$ fixed at $20,30,40$. §Empirical refers to the shared hyperparameter ranges for the three real-world datasets: ECG, Power Demand, and Astronomy, described in Sec. III.1.2.
C.2 Classification hyperparameter tuning
@@
For classification, the physical dimension $d$, maximum bond dimension $\chi_{\rm max}$ and learning rate $\eta$ were determined using the same hyperparameter tuning search algorithm in Appendix C.1 with 5-fold CV. The number of sweeps was fixed to $n_{\rm sweeps}=10$ across all experiments, and we used a scaled outlier robust sigmoid transformation to normalize the raw time-series data, as described in Sec. II.1. We evaluated $250$ MPS models for misclassification rate (i.e., 1 - accuracy) on each CV fold. The particular hyperparameter ranges we searched for each empirical dataset are summarized in Table 3.
Table 3: MPSTime imputation hyperparameter ranges for each dataset in Sec. III.3.
@@
$d$ [2, 15] [2, 15] [2, 15]
$\eta$ [0.001, 0.5]§ [0.001, 0.5]§ [0.001, 0.5]§
$\chi_{\rm max}$ [20, 40] [20, 40] [20, 40]
§$\eta$ was searched in logarithmic space to efficiently cover multiple orders of magnitude.